# iSeg: Interactive 3D Segmentation via Interactive Attention

### v1

@@ Abstract.
We present iSeg, a new interactive technique for segmenting 3D shapes. Previous works have focused mainly on leveraging pre-trained 2D foundation models for 3D segmentation based on text. However, text may be insufficient for accurately describing fine-grained spatial segmentations. Moreover, achieving a consistent 3D segmentation using a 2D model is challenging since occluded areas of the same semantic region may not be visible together from any 2D view. Thus, we design a segmentation method conditioned on fine user clicks, which operates entirely in 3D. Our system accepts user clicks directly on the shape’s surface, indicating the inclusion or exclusion of regions from the desired shape partition. To accommodate various click settings, we propose a novel interactive attention module capable of processing different numbers and types of clicks, enabling the training of a single unified interactive segmentation model. We apply iSeg to a myriad of shapes from different domains, demonstrating its versatility and faithfulness to the user’s specifications. Our project page is at https://threedle.github.io/iSeg/.
@@
ccs: Computing methodologies Shape representationsccs: Human-centered computing Interactive systems and toolsccs: Computing methodologies Image segmentation
[…]
@@ 1. Introduction
Interactive 3D segmentation is the ability to select fine-grained segments from a 3D shape that adhere to user inputs such as clicks. By contrast, traditional segmentation techniques do not utilize user inputs and instead rely on geometric features to delineate semantic parts (Hoffman and Richards, 1984; Dey and Zhao, 2004; Lien and Amato, 2007; Cornea et al., 2007; Shamir, 2008; Zheng et al., 2015).
In recent years, data-driven techniques have accelerated the 3D segmentation research by leveraging fully annotated 3D data (Milletari et al., 2016; Qi et al., 2017; Yi et al., 2017; Hanocka et al., 2019; Chen et al., 2019; Hu et al., 2022; Zhu et al., 2020; Lahav and Tal, 2020; Milano et al., 2020; Deng et al., 2020; Sharp et al., 2022; Sun et al., 2021). However, the reliance on a dataset and the scarcity of large-scale 3D datasets limits the network to a specific shape domain with a predefined set of parts.
Recent 3D segmentation techniques have circumvented the dependency on 3D data and pre-determined part definition by utilizing pretrained 2D foundation models to learn semantic co-segmentation (Ye et al., 2023) or text-driven segmentation (Kobayashi et al., 2022; Ha and Song, 2022; Decatur et al., 2023b; Abdelreheem et al., 2023b, a; Decatur et al., 2023a). However, text may not be able to accurately describe all fine-grained segmentations, such as the fourth leg of an octopus or a region corresponding to a particular point on the shape.
In this work, we present iSeg, a new data-driven interactive technique for 3D shape segmentation that generates tailored partitions of the shape according to user clicks. Given a shape represented as a triangular mesh, the user selects points on the mesh interactively to indicate a desired segmentation and iSeg predicts a region over the mesh surface that adheres to the clicked points. Our interactive interface can utilize positive and negative clicks, enabling additions and exclusions of areas from the segmented region, respectively (see for example Fig. 1).
We harness the power of a pretrained 2D foundation segmentation model (Kirillov et al., 2023) and distill its knowledge to 3D. However, segmenting a meaningful 3D region using a 2D model is very challenging, since occluded shape regions cannot be seen together from a single 2D view. Accordingly, we design an interactive segmentation system that operates entirely in 3D, where the user clicks and the inferred corresponding region are done over the shape surface directly, ensuring 3D consistency by construction. During training only, we project the 3D user clicks and the predicted segmentation to multiple 2D views to enjoy supervision from the powerful pretrained foundation model (Kirillov et al., 2023).
@@
Figure 3. Training of iSeg decoder. Our decoder takes the Mesh Feature Field (MFF) from iSeg encoder, along with the user input clicks, and generates a 3D segmentation map visualized in blue. We leverage a pre-trained 2D segmentation model (Kirillov et al., 2023) to supervise our training with 2D segmentation masks using rendered images of the shape and the 2D projection of the 3D clicks. Although iSeg is trained using noisy and inconsistent 2D segmentations, it is view-consistent by construction.
@@ 2. Related Work
Non-interactive 3D Segmentation. A large body of research has been focused on 3D segmentation using annotated datasets (Armeni et al., 2017; Hanocka et al., 2019; Hu et al., 2022; Lahav and Tal, 2020; Milano et al., 2020; Sharp et al., 2022; Yi et al., 2017). Such models demonstrate impressive performance at the cost of being restricted to the domain of the training data and the set of manually defined semantic labels. A partial solution to this limitation is utilizing unlabeled data, where common semantic elements are discovered by unsupervised learning (Chen et al., 2019; Deng et al., 2020; Hong et al., 2022; Sun et al., 2021; Zhu et al., 2020). Still, the segmentation is confined to the learned parts and is not easy to alter.
@@
Lifting 2D foundation models to 3D. The emergence of powerful 2D foundation models with a broad semantic understanding has propelled a surge of interest in distilling their knowledge and lifting it to a 3D representation (Kobayashi et al., 2022; Decatur et al., 2023b; Kundu et al., 2020; Zhang et al., 2022; Chen et al., 2023a; Ye et al., 2023; Kerr et al., 2023; Goel et al., 2023b; Abdelreheem et al., 2023b, a; Decatur et al., 2023a). Notably, (Kobayashi et al., 2022) and (Kerr et al., 2023) augmented the neural radiance scene representation (NeRF) (Mildenhall et al., 2020) with a volumetric feature filed. This approach enabled text-driven segmentation of objects within the scene, alleviating the need for a training dataset.
Similarly, we also lift the features of a 2D foundation model (Kirillov et al., 2023) into 3D. However, our method differs from previous work in several aspects. First, instead of using the implicit NeRF representation (Kobayashi et al., 2022; Kerr et al., 2023), our model operates directly on explicit 3D meshes, making it readily adaptable to 3D modeling workflows. Second, instead of learning an entire 3D feature volume, our mesh feature field lives only on the shape’s surface, simplifying the distillation process and making it more tractable. Third, rather than decoding the feature field with a simple correlation with the embedding the of semantic prompt, we learn a dedicated decoder in 3D to better exploit the semantic information embodied within our mesh feature field.
Interactive 3D Segmentation. Traditional interactive techniques have utilized heuristic smoothness priors and formulated the problem with a graph cut optimization objective (Boykov and Jolly, 2001; Rother et al., 2004; Sormann et al., 2006; Levin et al., 2007). More recently, several learning-based methods have been proposed for interactive segmentation (Goel et al., 2023a; Yue et al., 2023; Kontogianni et al., 2023). For example, (Kontogianni et al., 2023) segmented 3D point clouds based on user clicks. Unlike our work, they constructed a dataset for training their model, which limited its utility to parsing objects from a point cloud scene.
@@
Another approach taken by (Chen et al., 2023b) is to lift SAM’s features to a NeRF representation and use SAM’s decoder to obtain the segmentation masks. As in the first approach, applying the segmentation in 2D limits the method’s capabilities, since it cannot segment together occluded regions in 3D that are not visible concurrently in any 2D view. In contrast to existing works, our model and the user clicks are applied directly in 3D, simplifying the segmentation process and enabling the native segmentation of meaningful regions in 3D (as demonstrated in Sec. 3.2).
@@ 3. Method
We tackle the problem by proposing an interactive segmentation technique consisting of two parts: an encoder that maps vertex coordinates to a deep semantic vector, denoted as Mesh Feature Field (MFF), and a decoder that takes the MFF and the user clicks and predicts the corresponding mesh segment. The decoder contains an interactive attention layer supporting a variable number of clicks, which can be positive or negative, to increase or decrease the segmented region. Figure 3 presents an overview of the method.
@@ 3.1. Mesh Feature Field
The Mesh Feature Field (MFF) is a function $\phi\colon\mathbb{R}^{3}\rightarrow\mathbb{R}^{d}$ that embeds each mesh vertex into a deep feature vector $\phi(v_{i})=f_{i}$, where $d$ is the feature dimension. The MFF distills the semantic information from a pretrained 2D foundation model for image segmentation (Kirillov et al., 2023) and facilitates a 3D consistent feature representation for interactive segmentation of the mesh. To train the MFF, we render the high-dimensional vertex attributes differentiably and obtain the 2D projected features:
@@
where $\mathcal{R}$ is a differentiable renderer, $\theta$ is the viewing direction, and $w\times h$ are the spatial dimensions of the rasterized image.
The MFF is implemented as a multi-layer perceptron network, regarded as iSeg Encoder. To supervise its training, we render the mesh into a color image $I_{c}^{\theta}$ and pass it through the encoder $E_{2D}$ of the 2D foundation model (Kirillov et al., 2023) to obtain a reference feature map:
@@ 3.2. Interactive Attention Layer
Our interactive attention extends the scaled dot-product attention mechanism (Vaswani et al., 2017). We denote the encoded mesh features as $F\in\mathbb{R}^{n\times d}$. The features of the positively and negatively clicked vertices are marked as $F_{pos}\in\mathbb{R}^{n_{pos}\times d}$ and $F_{neg}\in\mathbb{R}^{n_{neg}\times d}$, respectively, where $n_{pos}$ and $n_{neg}$ are the number of clicks of each type. The interactive attention layer projects the mesh features to Queries and the features of the clicked points to Keys and Values:
(4) $\begin{split}Q=&FW^{Q}\in\mathbb{R}^{n\times d}\\ K_{\{pos,neg\}}=&F_{\{pos,neg\}}W^{K}_{\{pos,neg\}}\in\mathbb{R}^{n_{\{pos,neg\}}\times d}\\ V_{\{pos,neg\}}=&F_{\{pos,neg\}}W^{V}_{\{pos,neg\}}\in\mathbb{R}^{n_{\{pos,neg\}}\times d},\end{split}$
@@
Our attention mechanism condenses variable interactive user inputs into a fixed-length output. It learns the representation of positive and negative clicks, correlates the mesh vertices with them, and yields updated vertex features to enable the on-the-fly segmentation of the shape. Another benefit of our attention layer is that it is permutation invariant w.r.t. to the user clicks. In other words, it is independent of the sequential order of the point clicks and consistent in their joint influence on the shape partition. Moreover, the attention’s output $G$ is permutation equivariant w.r.t. to vertex order in $F$, a desired property for the mesh data structure.
[…]
@@ 3.3. Segmentation Prediction
To prepare data for our 3D decoder training, we simulate user clicks and generate probability masks from the 2D model (Kirillov et al., 2023). The data generation process includes 2 phases. First, we pick a subset of the mesh vertices uniformly, where each vertex is regarded as a single positive click. For each selected vertex, we generate random views, feed each one through the 2D foundation model, and get the reference segmentation mask. Then, for each view, we randomly sample another vertex visible within the viewing angle, which is set to be a positive or a negative click, and compute the updated segmentation by the 2D model. According to its type, we require the second click to increase or decrease the previous segmentation mask to obtain rich and diverse training data.
As seen in Fig. 3, the supervision signal is highly inconsistent. The 3D shape and the same clicks are interpreted differently by the 2D model, yielding strong variations in the 2D segmentation masks. Nonetheless, our method reveals a coherent underlying 3D segmentation function out of the noisy 2D measurements. Our decoder utilizes the robust distilled 3D vertex features, applies their interaction with the clicked points, and computes the region probability map directly in 3D. iSeg segmentaitons are view-consistent by construction, improving substantially over its training data. Furthermore, although trained with only 2D supervision, iSeg delineates meaningful regions in 3D that are not entirely visible in a single 2D projection (Sec. 3.2).
@@ 4. Experiments
We evaluate iSeg in a variety of aspects. First, we demonstrate the generality and fidelity of our method in Sections 4.1 and 4.2, respectively. Then, in Section 4.3, we showcase properties of iSeg, such as the stability of the segmentation for different clicks and its inherent 3D consistency. Finally, Section 4.4 presents the strong generalization power of iSeg in terms of the selected point, views of the click, and the number of clicks.
We apply our method to diverse meshes from different sources: COSEG (van Kaick et al., 2011), Turbo Squid (TurboSquid, 2021), Thingi10K (Zhou and Jacobson, 2016), Toys4k (Rehg, 2022), ModelNet (Wu et al., 2015), and ShapeNet (Chang et al., 2015). iSeg is highly robust to the shape properties. It operates on meshes with different numbers of vertices and various geometries, including thin, flat, and high-curvature surfaces.
iSeg is implemented in PyTorch (Paszke et al., 2017) and its training time varies according to the number of mesh vertices. For a mesh with $3000$ vertices, the optimization takes about 3 hours on a single Nvidia A40 GPU. Querying the model with input clicks takes about 0.7 seconds on that GPU. In our experiments, we used SAM ViT-H with an image size of $224\times 224$. To obtain fine-grained segmentations, we utilized the smallest scale mask from SAM for the projected clicked points. Additional details are in the supplementary.

### v2

@@ Abstract.
We present iSeg, a new interactive technique for segmenting 3D shapes. Previous works have focused mainly on leveraging pre-trained 2D foundation models for 3D segmentation based on text. However, text may be insufficient for accurately describing fine-grained spatial segmentations. Moreover, achieving a consistent 3D segmentation using a 2D model is highly challenging, since occluded areas of the same semantic region may not be visible together from any 2D view. Thus, we design a segmentation method conditioned on fine user clicks, which operates entirely in 3D. Our system accepts user clicks directly on the shape’s surface, indicating the inclusion or exclusion of regions from the desired shape partition. To accommodate various click settings, we propose a novel interactive attention module capable of processing different numbers and types of clicks, enabling the training of a single unified interactive segmentation model. We apply iSeg to a myriad of shapes from different domains, demonstrating its versatility and faithfulness to the user’s specifications. Our project page is at https://threedle.github.io/iSeg/.
@@
submissionid: 437journalyear: 2024copyright: rightsretainedconference: SIGGRAPH Asia 2024 Conference Papers; December 3–6, 2024; Tokyo, Japanbooktitle: SIGGRAPH Asia 2024 Conference Papers (SA Conference Papers ’24), December 3–6, 2024, Tokyo, Japandoi: 10.1145/3680528.3687605isbn: 979-8-4007-1131-2/24/12ccs: Computing methodologies Shape representationsccs: Human-centered computing Interactive systems and toolsccs: Computing methodologies Image segmentation
@@ 1. Introduction
Interactive 3D segmentation, the ability to select fine-grained segments from a 3D shape based on user inputs like clicks, is a fundamental problem in computer graphics with broad implications. In fields such as computer-aided design and 3D modeling, precise segment selection facilitates detailed model refinement. Moreover, in engineering, architecture, and medicine, fine-grained selection is indispensable for simulation and analysis, allowing for accurate assessment of structural integrity and behavior. While important, this problem poses significant challenges. How can we decipher the user intentions from such a minimal input as clicks? How do we handle diverse shapes with varying geometries and select specific and unique shape parts? In this work, we propose a method tailored to the shape at hand that selects regions adhering to the user clicks.
Traditional segmentation techniques do not utilize user inputs and instead rely on geometric features to delineate semantic parts (Hoffman and Richards, 1984; Dey and Zhao, 2004; Lien and Amato, 2007; Cornea et al., 2007; Shamir, 2008; Zheng et al., 2015). Recent data-driven techniques have further leveraged fully annotated 3D datasets and achieved ipressive 3D segmentation results (Milletari et al., 2016; Qi et al., 2017; Yi et al., 2017; Hanocka et al., 2019; Chen et al., 2019; Hu et al., 2022; Zhu et al., 2020; Milano et al., 2020; Deng et al., 2020; Sharp et al., 2022; Sun et al., 2021). However, the reliance on a dataset and the scarcity of large-scale 3D datasets limits the network to a specific shape domain with a predefined set of parts.
Current 3D segmentation methods have circumvented the dependency on 3D data and pre-determined part definition by utilizing pretrained 2D foundation models to learn semantic co-segmentation (Ye et al., 2023) or text-driven segmentation (Ha and Song, 2022; Decatur et al., 2023; Abdelreheem et al., 2023b, a; Liu et al., 2023; Decatur et al., 2024; Kim and Sung, 2024). Nonetheless, text may not be able to accurately describe all fine-grained segmentations, such as the fourth leg of an octopus or a region corresponding to a particular point on the shape.
@@
In this paper, we present iSeg, a new data-driven interactive technique for 3D shape segmentation that generates customized partitions of the shape according to user clicks. Given a shape represented as a triangular mesh, the user selects points on the mesh interactively to indicate a desired segmentation and iSeg predicts a region over the mesh surface that adheres to the clicked points. Our interactive interface can utilize positive and negative clicks, enabling additions and exclusions of areas from the segmented region, respectively (see Fig. 1).
We harness the power of a pretrained 2D foundation segmentation model (Kirillov et al., 2023) and distill its knowledge to 3D. However, segmenting a meaningful 3D region using a 2D model is very challenging, since occluded shape regions cannot be seen together from a single 2D view. Accordingly, we design an interactive segmentation system that operates entirely in 3D, where the user clicks and the inferred corresponding region are applied over the shape surface directly, ensuring 3D consistency by construction. During training only, we project the 3D user clicks and the predicted segmentation to multiple 2D views to enable supervision from the powerful pretrained foundation model (Kirillov et al., 2023).
@@
iSeg is optimized per mesh to capture its unique segments, without any ground-truth annotations. We train the model with only a small fraction of the mesh vertices, while the model successfully infers segmentations for other vertices not used during training. iSeg further generalizes beyond its training data and computes complete segments in 3D for clicks and regions occluded from each other.
@@ 2. Related Work
Figure 3. Training of the iSeg decoder. Our decoder takes the Mesh Feature Field (MFF) computed by the iSeg encoder, along with the user input clicks, and generates a 3D segmentation map visualized in blue. We leverage a pre-trained 2D segmentation model (Kirillov et al., 2023) to supervise our training with 2D segmentation masks using rendered images of the shape and the 2D projection of the 3D clicks. Although iSeg is trained using noisy and inconsistent 2D segmentations, it is view-consistent by construction.
2.1. Non-Interactive 3D Segmentation
A large body of research has been focused on 3D segmentation using annotated datasets (Armeni et al., 2017; Hanocka et al., 2019; Hu et al., 2022; Milano et al., 2020; Sharp et al., 2022; Yi et al., 2017; Kalogerakis et al., 2017). Such models demonstrate impressive performance at the cost of being restricted to the domain of the training data and the set of manually defined semantic labels. A partial solution to this limitation is utilizing unlabeled data, where common semantic elements are discovered by unsupervised learning (Chen et al., 2019; Deng et al., 2020; Hong et al., 2022; Sun et al., 2021; Zhu et al., 2020). Still, the segmentation is confined to the learned parts and is not easy to alter.
@@ 2.1. Non-Interactive 3D Segmentation
2.2. Lifting 2D Foundation Models to 3D
The emergence of powerful 2D foundation models with a broad semantic understanding has propelled a surge of interest in distilling their knowledge and lifting it to a 3D representation (Decatur et al., 2023; Kundu et al., 2020; Zhang et al., 2022; Chen et al., 2023a; Abdelreheem et al., 2023b, a; Peng et al., 2023; Fan et al., 2023; Umam et al., 2024; Yin et al., 2024; Decatur et al., 2024). Notably, several researchers (Tschernezki et al., 2022; Kobayashi et al., 2022; Kerr et al., 2023) augmented the neural radiance scene representation (NeRF) (Mildenhall et al., 2020) with a volumetric feature filed. This approach enabled text-driven segmentation of objects within the scene, alleviating the need for a training dataset.
Similarly, we lift the features of a 2D foundation model (Kirillov et al., 2023) into 3D. However, instead of using the implicit NeRF representation (Ye et al., 2023; Kerr et al., 2023), our model operates directly on explicit 3D meshes, making it readily adaptable to 3D modeling workflows. Moreover, rather than decoding the feature field by a simple correlation with the embedding of the semantic prompt (Fan et al., 2023; Peng et al., 2023), we learn a dedicated decoder in 3D to exploit the semantic information embodied within our mesh feature field better.
2.3. Interactive 3D Segmentation
Traditional interactive techniques have utilized heuristic smoothness priors and formulated the problem with a graph cut optimization objective (Boykov and Jolly, 2001; Rother et al., 2004; Sormann et al., 2006). More recently, several learning-based methods have been proposed for interactive segmentation (Ren et al., 2022; Goel et al., 2023; Yue et al., 2024; Kontogianni et al., 2023; Mirzaei et al., 2023; Ying et al., 2024). For example, (Kontogianni et al., 2023) segmented 3D point clouds based on user clicks. Unlike our work, they constructed a dataset for training their model, which limited its utility to parsing objects from a scene.
@@ 2.3. Interactive 3D Segmentation
Another approach taken by (Chen et al., 2023b) is to lift SAM’s features to a NeRF representation and use SAM’s decoder to obtain the segmentation masks. As in the first approach, applying the segmentation in 2D limits the method’s capabilities, since it cannot segment together occluded regions in 3D that are not visible concurrently in any 2D view. In contrast to existing works, our model and the user clicks are applied directly in 3D, simplifying the segmentation process and enabling the native segmentation of meaningful regions in 3D (as demonstrated in Fig. 5).
@@ 3. Method
We tackle the problem by proposing an interactive segmentation technique consisting of two parts: an encoder that maps vertex coordinates to a deep semantic vector and a decoder that takes the vertex features and the user clicks and predicts the corresponding mesh segment. The decoder contains an interactive attention layer supporting a variable number of clicks, which can be positive or negative, to increase or decrease the segmented region. Fig. 3 presents an overview of the method.
@@ 3.1. Mesh Feature Field
Our encoder learns a function $\phi\colon\mathbb{R}^{3}\rightarrow\mathbb{R}^{d}$ that embeds each mesh vertex into a deep feature vector $\phi(v_{i})=f_{i}$, where $d$ is the feature dimension. The collection of mesh vertex features is denoted as $F\in\mathbb{R}^{n\times d}$ and regarded as the Mesh Feature Field (MFF).
The encoder distills the semantic information from a pretrained 2D foundation model for image segmentation (Kirillov et al., 2023) and facilitates a 3D consistent feature representation for interactive segmentation of the mesh. To train the encoder, we render the high-dimensional vertex attributes differentiably and obtain the 2D projected features:
@@
where $\mathcal{R}$ is a differentiable renderer, $\theta$ is the viewing direction, $f$ represents the visible vertices in the view, and $w\times h$ are the spatial dimensions of the rasterized image.
The encoder is implemented as a multi-layer perceptron network. To supervise its training, we render the mesh into a color image $I_{c}^{\theta}$ and pass it through the encoder $E_{2D}$ of the 2D foundation model (Kirillov et al., 2023) to obtain a reference feature map:
@@ 3.2. Interactive Attention Layer
Our interactive attention extends the scaled dot-product attention mechanism (Vaswani et al., 2017). The features of the positively and negatively clicked vertices are marked as $F_{pos}\in\mathbb{R}^{n_{pos}\times d}$ and $F_{neg}\in\mathbb{R}^{n_{neg}\times d}$, respectively, where $n_{pos}$ and $n_{neg}$ are the number of clicks of each type. The interactive attention layer projects the mesh features $F$ to Queries and the features of the clicked points to Keys and Values:
(4) $\begin{split}Q=&\>FW^{Q}\in\mathbb{R}^{n\times d}\\ K_{\{pos,neg\}}=&\>F_{\{pos,neg\}}W^{K}_{\{pos,neg\}}\in\mathbb{R}^{n_{\{pos,neg\}}\times d}\\ V_{\{pos,neg\}}=&\>F_{\{pos,neg\}}W^{V}_{\{pos,neg\}}\in\mathbb{R}^{n_{\{pos,neg\}}\times d},\end{split}$
@@
Our attention mechanism condenses variable interactive user inputs into a fixed-length output. It learns the representation of positive and negative clicks, correlates the mesh vertices with them, and yields updated vertex features to enable the on-the-fly segmentation of the shape. Another benefit of our attention layer is that it is permutation invariant w.r.t. the user clicks. In other words, it is independent of the sequential order of the point clicks and consistent in their joint influence on the shape partition. Moreover, the attention’s output $G$ is permutation equivariant w.r.t. the vertex order in $F$, a desired property for the mesh data structure.
@@ 3.3. Segmentation Prediction
To prepare data for our 3D decoder training, we simulate user clicks and generate masks from the 2D model (Kirillov et al., 2023). The data generation process includes two phases. First, we pick a small training subset of 3% of the mesh vertices well distributed over the shape using Farthest Point Sampling (Eldar et al., 1997), where each vertex is regarded as a single positive click. For each vertex, we generate random views, feed each one through the 2D foundation model, and get the reference segmentation mask. Then, for each view, we sample another training vertex visible within the viewing angle, which is set to be a positive or a negative click, and compute the updated segmentation by the 2D model. According to its type, we require the second click to increase or decrease the previous segmentation mask to obtain rich and diverse training data.
As seen in Fig. 3, the supervision signal is highly inconsistent. The 3D shape and the same clicks are interpreted differently by the 2D model, yielding strong variations in the 2D segmentation masks. Nevertheless, our method reveals a coherent underlying 3D segmentation function out of the noisy 2D measurements. Our decoder utilizes the robust distilled 3D vertex features, applies their interaction with the clicked points, and computes the region probability map directly in 3D. iSeg segmentaitons are view-consistent by construction, improving substantially over its training data. Furthermore, although trained with only 2D supervision, iSeg delineates meaningful regions in 3D that are not entirely visible in a single 2D projection (Fig. 5).
@@ 4. Experiments
We evaluate iSeg in a variety of aspects. First, we demonstrate the generality and fidelity of our method in Secs. 4.1 and 4.2, respectively. Then, in Sec. 4.3, we showcase the generic feature information captured by iSeg. Finally, Sec. 4.4 presents the strong generalization power of iSeg in terms of the selected point, views of the click, and the number of clicks.
We apply our method to diverse meshes from different sources: COSEG (van Kaick et al., 2011), Turbo Squid (TurboSquid, 2021), Thingi10K (Zhou and Jacobson, 2016), Toys4k (Rehg, 2022), SCAPE (Anguelov et al., 2005), SHREC ’19 (Melzi et al., 2019), ModelNet (Wu et al., 2015), ShapeNet (Chang et al., 2015), and PartNet (Mo et al., 2019). iSeg is highly robust to the shape properties. It operates on meshes with different numbers of vertices and various geometries, including thin, flat, and high-curvature surfaces.
iSeg is implemented in PyTorch (Paszke et al., 2017) and its training time varies according to the number of mesh vertices. For a mesh with $3000$ vertices, the optimization takes about 3 hours on a single Nvidia A40 GPU. Training the model is a one-time offline phase. Once trained, querying the model with input clicks takes only about 0.7 seconds, which allows fast interactive interaction with the shape. In our experiments, we used SAM ViT-H with an image size of $224\times 224$. To obtain fine-grained segmentations, we utilized the smallest scale mask from SAM for the projected clicked points. Additional details are provided in the supplementary.